OpenStudy (anonymous):

(3n-6)n ------- = -2(n-2) A. -n B. -3n/4 C. -n/2 D. -3n/2

OpenStudy (anonymous):

The answer is D however idk how to get that

OpenStudy (anonymous):

\[\frac{ (3n-6)n }{ -2(n-2) }=\frac{ 3n^2-6n }{ -2(n-2) }\]

OpenStudy (anonymous):

How???

OpenStudy (anonymous):

\[=\frac{ 3n(n-2) }{ -2(n-2) }\]

OpenStudy (anonymous):

what u get when u multiply 3n-6 with n?

OpenStudy (anonymous):

3n squared

OpenStudy (anonymous):

and 6n

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

YAY

OpenStudy (anonymous):

\[=\frac{ 3n^2-6n }{ -2(n-2) }\] now factorise 3n for the numerator only

OpenStudy (anonymous):

you will get \[=\frac{ 3n(n-2) }{ -2(n-2) }\]

OpenStudy (anonymous):

and you did that how again

OpenStudy (anonymous):

facotrise 3n for the numerator only

OpenStudy (anonymous):

factorise*

OpenStudy (anonymous):

ik but how what did you use to do that

OpenStudy (anonymous):

Sorry I get really confused on these type of questions

OpenStudy (anonymous):

Thinking, of how to explain it to me?

geerky42 (geerky42):

Well, sometime we used GCF to see what's GCF in terms inside of parenthesis, then factor that out. \(\text{GCF}(3n^2,6n) = 3n\)\ so factor \(3n\) out.

OpenStudy (anonymous):

I don't think I've learn GCF which is making me wonder why are they giving me this to study for my end of year test, oh dear

OpenStudy (anonymous):

so basically I don't know how to use GCF so can someone explain that to me